### v2

# mpdframe frametitlebackgroundcolor =black!15, frametitlerule =true, roundcorner =1pt, middlelinewidth =1pt, skipabove =nermargin =0.1cm, innerleftmargin =0.1cm, innerrightmargin =0.1cm, innertopmargin =0.1cm, innerbottommargin =0.1cm

Abstract
N-Version Programming is a well-known methodology for developing fault-tolerant systems. It achieves fault detection and correction at runtime by adding diverse redundancy into programs, minimizing fault mode overlap between redundant program variants. In this work, we propose the automated generation of program variants using large language models. We design, develop and evaluate Galápagos: a tool for generating program variants using LLMs, validating their correctness and equivalence, and using them to assemble N-Version binaries. We evaluate Galápagos by creating N-Version components of real-world C code. Our original results show that Galápagos can produce program variants that are proven to be functionally equivalent, even when the variants are written in a different programming language. Our systematic diversity measurement indicates that functionally equivalent variants produced by Galápagos, are statically different after compilation, and present diverging internal behavior at runtime. We demonstrate that the variants produced by Galápagos can protect C code against real miscompilation bugs which affect the Clang compiler. Overall, our paper shows that producing N-Version software can be drastically automated by advanced usage of practical formal verification and generative language models.
I Introduction
N-Version programming is a software engineering approach to reliability based on crafting software diversity with different teams [1]. It is predominantly used to enhance the reliability of mission-critical systems [2]. This is achieved by building diverse and equivalent software components, with the objective that these would not fail all at once, under the same conditions, hence making the whole system more robust [2].
The largest challenge of N-Version programming relates to the development and operation costs of the approach [3]. There are high costs induced by the manual development and verification of multiple program versions instead of a single one, typically linear in the number of versions. Another important additional cost is the development of additional software to orchestrate the multiple versions (the N-version harness).
Large Language Models (LLMs) are powerful tools able to carry out complex source code-related tasks, such as code completion [4], automatic program repair [5], and code summarization [6]. Our key intuition is that they are able to efficiently automate N-Version programming. Given a reference program, we can leverage the creative and search power of LLMs to automatically produce program variants. We anticipate these variants to be diverse, due to the great volume of data used in LLM training, apt to being combined in an N-Version system. In other words, we want to harness the LLM’s ability to synthesize program variants, and instead of having multiple teams producing multiple versions at a high costs, having the LLM producing them at virtually no cost.
In this paper, we design, implement and evaluate Galápagos, a tool for automated, verified N-Version programming. Galápagos is made up of three passes, which use off-the-shelf LLMs and harness program equivalence verification to provide strong guarantees. Galápagos generates variants of a reference function and then assembles N-Version implementations of that function. We implement Galápagos in the scope of C libraries, with support for function variant generation as well as N-Version in same-language (C-to-C), and cross-language (C-to-Go).
1//main.c
2int printf(const char *, ...);
3static int a = -3, b;
4static char c;
5int d;
6int e(int f, int g) {
7 if (f - g < 10000)
8 return f;
9 return f + 1 % -f;
10}
11int main() {
12 int *h[] = {&a, &a};
13 for (; c <= 37; ++c) {
14 int *i = &b;
15 *i |= e(a, 8) + d;
16 }
[…]
18}
\$ clang -v
clang version 17.0.0
Target: x86_64-unknown-linux-gnu
\$ clang -O0 main.c; ./a.out
-3
\$ clang -Os main.c; ./a.out
-1
Figure 1: C code which triggers a miscompilation on clang v17.0.0, when using the -Os flag. After running the compiled binary, it prints -1. The correct result is -3. Galápagos is able to detect and mitigate this miscompilation fault.
We conduct a systematic evaluation of Galápagos to quantify the correctness, diversity, and usefulness of both the function variants and the N-Version components. First, we assemble a dataset of 30 real-world C functions, extracted from 6 notable cryptography and multimedia open source projects. We run Galápagos to transform each function in the dataset into an N-Version component. Second, we measure the diversity of the variants, both statically and dynamically. Statically, we measure the uniqueness of the resulting machine code after compiling the variants, using different optimization configurations. Dynamically, we execute each unique variant, and record the CPU instructions they utilize. This allows us to quantify the variation in execution paths, which is known to be important against certain types of exploits such as side-channel attacks [7]. Third, we rigorously test the assembled N-Version implementations in the context of miscompilation bugs [8]. It clearly demonstrates Galápagos’s ability to enhance robustness wrt. our target fault model.
Our experiments demonstrate several key findings: First, Galápagos uses LLMs to generate code variants that are provably equivalent, even when these variants are written in different programming languages. During our experiments, 234 equivalent variants were found from a dataset of 30 reference programs. Second, Galápagos is capable of producing and identifying code variants that exhibit diversity both in their code (statically) and during execution (dynamically). From the dataset of 30 programs, 126 unique variants were verified as different, even after all compilation and optimization passes. Last but not least, these diverse code variants can be utilized to strengthen critical sections of software via automated N-Version programming: we demonstrate Galápagos’ capability to mitigate real-world, reported miscompilation bugs in the Clang compiler.
In summary, our contributions are
  • The design and implementation of Galápagos, a tool for automated and verified N-Version programming using LLMs. As far as we are aware, this is the first work to realize an automatic N-Version programming framework with formal guarantees. Galápagos is a major contribution towards reducing the high costs of N-Version programming.
  • A large scale suite of experiments to evaluate the correctness, value, and practicality of automated N-Version programming with Galápagos. The experiments consist of transforming real-world C code from notable open-source libraries into an N-Version counterpart.
  • A publicly available, open-source repository for experimental reproduction and future research on automated N-Version programming, at  https://github.com/ASSERT-KTH/Galapagos/
II Background
In this section, we introduce the two areas in which our work is rooted: N-Version programming and neural machine translation.
II-A N-Version Programming
N-Version programming is a software development approach, which consists of creating $N$ implementations or versions of a specific program [1]. The core idea behind this approach is that when these versions are simultaneously executed, errors can be timely detected and mitigated by comparing their outputs. Ideally, the difference in implementations between versions is maximal, such that any coincidental errors are avoided. [9]. While originally devised as a fault-tolerance mechanism, N-Version programming has been adapted to enhance other specific properties of software, such as availability [10], reliability [11], performance [12], or security [13, 14]
However, the enhancements offered by N-Version programming come with an attached trade-off, as it introduces many challenges throughout the software development lifecycle. These challenges include increased maintenance overhead, increased compute and memory use, or interoperability issues [15, 10]. Addressing these challenges requires additional effort and careful coordination across engineering teams.
Figure 2: Galapagos pipeline for automated, verified N-Version programming.
An essential challenge is the increased cost of development, given that the time and resources needed to develop the versions increase at least linearly with N. To address this challenge, automating the process for creating new versions is a known and well-studied approach [16, 7, 17, 18]. In this paper, we contribute to this field of automatic synthesis of diverse program versions.
II-B Code Translation with LLMs
Neural machine translation (NMT) for source code refers to the application of machine learning techniques for the automated translation between programming languages [19]. It aims to overcome the challenges associated with traditional rule-based or statistical translation methods by leveraging the power of neural networks to capture and model complex patterns inherent in source code [20]. The core principle behind NMT for source code involves training a neural network model on large datasets consisting of source code snippets in different programming languages and their corresponding translations [21]. During the training phase, the model learns to encode the source code syntax and semantics into a continuous representation, enabling it to generate accurate translations. Subsequently, the trained model can be deployed to translate code snippets from one programming language to another automatically. NMT is tied to recent developments in large language models, which have proven to be efficient in performing translation tasks [22].
III N-Version programming with LLMs
In this section, we present Galápagos, a tool that leverages large language models to harden software against miscompilation errors via automated N-Version programming . Galápagos works at the source-code level for generation, and at the intermediate representation (IR) level for verification. At the source-code level, it uses an LLM to produce new, diversified variants of program functions. At the intermediate representation level, these variants are verified to be semantically equivalent and to be finally assembled into N-Version functions. Because the variants are formally verified to be equivalent, Galápagos is able to provide strong guarantees about the behavior final N-Version assembly, irrespective of the introduced software diversity.
III-A Fault Model
In this paper, we propose a novel technique to mitigate the class of faults called miscompilation. Miscompilations are caused by bugs in compilers, resulting in output machine code not matching the behavior defined in the source code. These faults are very hard to identify, since they are silent, and generally can only be detected indirectly after analyzing misbehaving applications [23].
Miscompilations are present in all major compilers [24]. As of July 2024, Clang and GCC have 340 and 47 unresolved reports of miscompilation bugs, respectively. As an example, consider the code from an open LLVM bug report in Figure 1. After compilation using clang v17.0.0, with the -Os optimization flag, the result is a binary which produces incorrect output: -1. The compilation process exits “successfully” with code 0, and offers no warnings or error messages which would hint that something has gone wrong. Hence, there is an error in one of Clang’s optimization passes. This kind of fault can cause major downstream failures [25].
Let us abstract the compilation process as the following composition of functions:
$P=B_{a}(M(F_{l}(S)))$ (1)
In Equation 1, $S$ is a program’s source code; $F_{l}$ is the front-end stage, which is specific to the language $l$, and transforms $S$ to an intermediate representation (IR); $M$ is the middle-end phase, i.e. a set of optimization passes performed on the IR; $B_{a}$ is the back-end stage, which transforms the IR into assembly code targeted towards a specific architecture ${a}$; and $P$ is the resulting executable program.
In this work, we want to harden programs generated by the compilers’ middle-end $M$ and back-end $B_{a}$ stages, i.e. any operations performed on the intermediate representation which result in $P$ not behaving as specified by $S$.
III-B Overview
Galápagos is a three-pass pipeline, as illustrated in Figure 2: Diversification, Validation, and Harnessing.
The pipeline’s input is a function’s source code, which is part of a larger application or library. For instance, a function could compute a cryptographic hash. Hereafter, we refer to this function as the reference function or simply the reference. The reference defines the expected functionality of the function on the whole input domain. In other words, the reference formally specifies the input/output behavior that all variants should conform to.
To create an N-Version implementation of the reference function, it is first processed in the Diversification pass (subsection III-C) of Galápagos. Here, an LLM is prompted to automatically generate different variants of the reference. Second, the collection of generated variants is passed to the Validation pass (subsection III-D), which filters out non-viable variants produced by the Diversification pass. The Validation pass proceeds through a sequence of validation steps: (1) Galápagos compiles each variant and produces its corresponding IR code, variants that cannot be compiled are filtered out as non-viable; (2) The IR code of all variants that compile is tested, and sent to a formal equivalence checker to compare their functionality against the reference. The variants that pass the tests and the equivalence check are then forwarded to the next pass of Galápagos. This ensures that after this pass, the ensuing variants have formal guarantees that the original functionality is preserved, and expected diverse internal behavior per the LLM instructions. Third, the Harnessing pass (subsection III-E) uses the resulting variants to assemble an executable, where the original function is replaced by an N-Version implementation of that function. The specific LLM and equivalence checker are external and configurable, which allows Galápagos to use any off-the-shelf solutions.
Galápagos generates a binary that contains an N-Version implementation of the reference function. This binary protects against miscompilation faults, thanks to two properties: (1) miscompilations are very sensitive to the input source code, e.g. they might be related to a specific keyword [26]; consequently, deriving program variants from diverse-but-equivalent source code is less likely to trigger miscompilations in every diversified instance, and; (2) miscompilations are compiler-specific [8], meaning that it is improbable that variants produced by different compilers (see subsection III-C) share the same miscompilation error.
III-C Diversification Pass
The goal of this step is to systematically create function variants given a reference function.
As shown in Figure 2, the Diversification pass is managed by a controller module. The controller is responsible for reading configuration, interfacing with LLM APIs, and processing responses. The invocation of the LLM is configured by the following parameters: the LLM API where the diversification request will be sent to; the reference function’s source code; the number of variants to be generated; and a pair of input and output languages. The input language is the one of the reference function. The output language can be the same one as the reference function’s (e.g. diversifying C-to-C), or a different language (e.g. diversifying C-to-Go).
In Galápagos, the LLM prompt structure is as follows:
  • A natural language description of the reference function: The following code is a reference implementation of a function in <INPUT>.
  • The reference function.
  • A description of the task: Create <NUMBER> substitute implementation(s) of the function [in the <OUTPUT> language], which are different but equivalent. It should be possible to directly replace the function with any substitute, and it should provide the same functionality.
  • Additional remarks: Do not output any other text apart from code. Do not create auxiliary or helper functions. Maintain the original function’s signature.
The <NUMBER>, <INPUT>, and <OUTPUT> strings are replaced according to the configuration, and the text in square brackets is omitted if the input and output languages are the same.
Figure 3: Galápagos’ diversification pass example. On the left box, the Galápagos prompt; on the right box, the response from the LLM.
Figure 3 shows an example of the diversification pass, including the resulting prompt after configuration, and resulting function variants. In this example, the configuration is set to generate three Go variants from a C reference function. In the response, we can observe that while the first variant is a direct translation of the reference function, the second and third variants are dissimilar, using extra operations or different control flow constructs.
As shown in the previous example, the diversification pass can be performed in a cross-language configuration. This is possible because the Validation and Harnessing passes of Galápagos are performed at the IR level. Hence, the source languages for the reference and the variants can be different, as long as there exist front ends that can compile the languages to the same IR.
This diversification approach presents two main challenges. First, LLM outputs are not always correct [27]. We need a mechanism to guarantee that the produced variants are equivalent. This concern is addressed in the Validation pass (subsection III-D) Second, when diversifying the reference to different programming languages, the IR code resulting from compiling the variants might not be compatible. We mitigate this concern in the Harnessing pass (subsection III-E).
To sum up, the Diversification pass of Galápagos crafts an LLM prompt to produce variants of the reference function either in the same language or in a different language that can be transformed into the same IR. The resulting variants are forwarded to Galápagos’ next pass.
III-D Validation Pass
The goal of the Validation pass is to guarantee that the functionality of the function variants produced by the Diversification pass is equivalent to the reference.
In contrast to previous work where automatic diversification is achieved by performing transformations known to preserve equivalence  [26, 16, 7], Galápagos leverages LLMs, and thus the variants are not guaranteed to be correct by construction [28, 29]. To address this issue, Galápagos performs four checks on each variant: compilation in isolation, compilation within containing project, test suite success, and equivalence check.
Compilation success in isolation: In this check, Galápagos inserts the variant into a file which only contains an entry point, e.g., a main function, that calls the variant with arbitrary parameters. Then we attempt to compile this file with the corresponding language’s toolchain. The variant is passed to the next step if compilation completes without any error.
Compilation success within the corresponding project: In this check, Galápagos replaces the reference function’s IR inside the global project with the variant’s IR. The variant is filtered out if the build pipeline is unsuccessful. Reasons for failure at this step are related to IR incompatibility, e.g. mismatching types.
Test suite success: In this check, Galápagos executes the test suite against each binary that includes a valid variant created in the previous step. The variant is filtered out if any test from the suite does not pass. At this step, a test case that fails is direct evidence of a functionality difference between the original and the variant.
Equivalence check: In this final check, Galápagos calls a configurable off-the-shelf formal equivalence checking tool, such as alive2 [30], or Rust’s Kani [31]. This uncouples Galápagos from any specific IR language and allows the use of any state-of-the-art solution for verification. It performs the equivalence check with the reference and variant IR codes as parameters. The variant is filtered out if the tool proves the variant to be non-equal, or if it fails to prove equivalence within a given time and memory budget. Reasons for failure at this step are direct evidence of non-equivalence, as the variant deviates from the specification’s functionality. This is the most resource-consuming check, so it is executed last. All function variants that pass the equivalence check are considered correct and can be forwarded to the next pass for assembly. Addressing the accuracy of the equivalence checking tool is out of the scope of this work. We point that equivalence check false negatives would not introduce errors to the system, as the variant would be filtered out. Equivalence check false positives are more challenging, as non-equivalent variants would be accepted in the N-version assembly, Yet, their divergent output would not propagate silent corruptions, but instead would trigger a program termination (see  subsection III-E).
To sum up, the Validation pass filters candidate variants, eventually keeping only the variants for which it can provide formal correctness guarantees. Galápagos blends the radiating diversity powered by the LLM’s creativity with strict guarantees of formal program equivalence checking.
III-E Harnessing Pass
Per N-Version programming, the function variants need to be assembled into a single unit before execution. In Galápagos, the assembling is done at the intermediate representation level.
The previous passes of Galápagos aim at maximizing diversity. They leverage the diversity of source languages and blend multiple toolchains. Therefore, the variants’ IR contain subtle differences, which is at the core of the software diversity we are seeking. Yet, these differences need to be dealt with when assembling N-version functions. To handle this case, Galápagos supports non-trivial IR transformations of the variants during the harnessing pass. Specifically, Galápagos implements Go-to-C transformations related to function signatures and replacing language-specific components: (1) the LLVM function’s nest self-reference is removed; (2) the function’s parameter names are normalized; and (3) IR calls to the Go "runtime" module are removed  [^1] https://github.com/ASSERT-KTH/Galapagos/blob/main/linker/lib/go/go.cpp.
Figure 4 shows an example where two equivalent variants of a function which adds two integers are harnessed into an N-Version function. This example is written in the LLVM intermediate representation. The variants’ code is shown in the boxes on the left-hand side, and the resulting N-Version function is shown in the box on the right-hand side. In the N-Version form, the code from the single variants is inserted in the corresponding file as a function definition. These definitions are named after the original function plus a prefix. Galápagos then generates a wrapper function named after the original reference function. This function invokes all the different variants and compares their outputs. Galápagos also transforms the original application to redirect all existing invocations of the reference to the wrapper function. Finally, the binary is produced from the resulting IR code.
Figure 4: Galápagos’s harnessing pass example, in LLVM IR.
For N-Version programs, the output from the N functions can be selected in different ways. In Galápagos, we support N-of-N selection [32], i.e. we compare the return values of each variant against each other and return a value only if all the variants agree, otherwise Galápagos forcefully terminates the execution.
III-F Class of Functions Considered for Automated N-Version
Galápagos transforms functions into N-Version functions, with equivalence guarantees. To achieve this challenging task, we must assume a specific class of functions. Namely, these functions must be pure since all versions must be executed simultaneously, per the practice of N-Version programming. Only with pure functions can we ensure that multiple executions of the function will not result in accumulated, undesired side effects.
This scope for Galápagos is relevant, as pure functions represent a significant subset of all functions in software [33, 34]. Figure 5 shows a relevant example from the cryptography library libsodium that can be divsersified. Furthermore, we argue that in practice, critical sections of code that have strong reliability requirements can be refactored into pure functions, making hardening via Galápagos and N-Version programming possible.
1static int
2_crypto_scalarmult_ed25519_is_inf(const unsigned char s[32])
3{
4 unsigned char c;
5 unsigned int i;
6
7 c = s[0] ^ 0x01;
8 for (i = 1; i < 31; i++) {
9 c |= s[i];
10 }
11 c |= s[31] & 0x7f;
12
13 return ((((unsigned int) c) - 1U) >> 8) & 1;
14}
Figure 5: Pure function example from the cryptography library libsodium, which meets the criteria for automated N-Version programming with Galápagos.
III-G Implementation
We implement the Diversification and Validation passes of Galápagos in Python, and make use of OpenAI’s GPT-4o API. The Harnessing pass is implemented in C++ in order to make use of LLVM’s code manipulation libraries. For formal equivalence checks, we rely on the alive-tv [30] tool that checks equivalence at the LLVM IR level.
Galápagos supports C as input language, and C and Go as output languages. Support can be extended to any language pair, as long as these can be compiled to a common intermediate representation.
The code can be found in our open source repository at https://github.com/ASSERT-KTH/Galapagos.
IV Experimental Methodology
In this section, we introduce the research questions that structure our empirical validation of Galápagos.
IV-A Research Questions
To systematically evaluate the effectiveness of Galápagos, we define the following research questions:
  1. RQ1
    To what extent is a large language model able to produce functionally equivalent variants of programs at the source-code level?
    The variants generated by an LLM come with no guarantee. We aim to compute what percentage of these variants can be proven to perform the same task as the original function under diversification. To answer this question, we take a representative sample of functions from relevant C applications and libraries, and run them through Galápagos’ pipeline. Then, we analyze the distribution of these variants in terms of the different correctness filters of Galápagos’ validation passes presented in subsection III-D.
  2. RQ2
    To what extent do the program variants produced by a large language model exhibit diverse observable internal behaviors?
    The variants generated by an LLM can be different at the source code or IR code. However, program variants stored on disk do not increase reliability per se. N-Version design increases reliability when variants exhibit observable differences. For this, we compute how much of the variants’ diversity is preserved after all compilation and optimization stages. Furthermore, we compare how much CPU instructions vary between different function variants.
  3. RQ3
    To what extent are equivalent variants produced by a large language model useful to harden programs against the considered fault model?
    We evaluate the effectiveness of N-Version functions per our fault model, miscompilation bugs (see subsection III-A). We curate publicly documented, reproducible miscompilation bugs, and determine whether LLM-generated variants are useful in detecting and preventing them.
  4. RQ4
    What is the performance overhead introduced by the N-Version functions generated by Galápagos?
    N-Version functions with high execution overhead might be impractical. Therefore, we measure the resulting performance overhead introduced by Galápagos. To do this, we run the original and N-Version functions and compare their average execution times.
IV-B Dataset of Functions
We select C functions from six open-source projects: alsa-lib, FFmpeg, libgcrypt, liboqs, libsodium, openssl. Those projects are all real-world, and have differing sizes, stages of maturity, and adoption.  Table I shows the selected projects, their versions, and related statistics. The number of commits for the selected projects’ repositories ranges between 1 458 and 111 662. Likewise, the number of C files and declared functions goes from 349 to 4 160, and 3 683 to 110 532 respectively.
Project Version # Commits # Files # Functions
alsa-lib v1.2.11* 4 445 400 4 004
FFmpeg n6.1-dev 111 662 4 160 110 532
libgcrypt 1.10-base* 3 419 588 5 616
liboqs 0.10.0* 1 458 3 625 9 594
libsodium 1.0.18 4 204 349 3 683
openssl 3.0.0-beta2 32 959 4 004 94 984
TABLE I: List of real-world C projects used for experimental validation,. The versions marked with (*) represent the oldest named version before the commit used in this dataset.
From each of these projects, we select five functions that meet the following properties. (1) It is a pure function, and; (2) It does not call other functions. These properties ensure that the functions are supported by the semantic equivalence tool used in Galápagos’ verification step.
Project Function Name Size (LOC) # Refs.
alaw_to_s16 15 1
iec958_parity 13 1
add 6 3
ulaw_to_s16 15 1
alsa-lib val_seg 15 2
flac_get_max_frame_size 19 2
mix 6 43
vp5_adjust* 14 2
weight 10 2
FFmpeg int_sin 17 5
barrett_reduce 7 2
ctz 12 1
int16_t_negative_mask 7 2
int16_t_nonzero_mask 7 5
libgcrypt montgomery_reduce 7 2
fpr_half 11 132
fpr_lt 29 80
modp_montymul 9 1 220
modp_norm 2 80
liboqs int16_nonzero_mask 6 16
b64_byte_to_char 6 4
b64_byte_to_urlsafe_char 6 4
b64_char_to_byte 9 2
b64_urlsafe_char_to_byte 9 2
libsodium fBlaMka 5 28
icbrt64 15 1
BitDeinterleave 34 1
BitInterleave 34 1
_booth_recode_w5 10 3
openssl _booth_recode_w7 10 2
TABLE II: List of 30 candidate real-world C functions for diversification.
Table II describes the set of functions we selected, which covers a diverse range of functionalities, e.g. arithmetic, type conversion, and cryptographic operations. It also shows each function’s size in lines of code, and the number of times that it is statically referenced within the project. The size of the functions by lines of code ranges from 2 to 34 with a median of 10. The number of times that each function is called within the same project ranges from 1 to 1 220 with a median of 2. To sum up, this dataset is exclusively composed of real-world C code that is used daily by millions of users.
Tag Version Description Issue # Status
M1 15.0.0 Wrong code with -O1 58765 Unresolved
M2 17.0.0 Wrong code with -Os 68871 Unresolved
M3 18.0.0 Wrong code with -march=znver4 -flto -O3 80494 Unresolved
TABLE III: List of known miscompilation bugs selected for mitigation through Galápagos.
IV-C Methodology for RQ1: Assessing Functions Diversified by LLMs
As part of its Diversification pass, Galápagos generates an arbitrary number of raw variants. However, there are mandatory properties which every variant needs to comply with for N-Version design: compilability, test suite execution success, and formal equivalence assessment. To have a detailed view of the distribution of equivalent and non-equivalent variants, we measure the proportion of variants which fail at each filter and the reasons behind the failures. Ultimately, the effectiveness of LLM-based diversification is measured by the proportion of variants which pass all filters and are proven equivalent.
Importantly, we diversify the dataset of C functions (subsection IV-B) with two configurations same-language (C-to-C) and cross-language (C-to-Go). We produce 10 variants for each configuration, i.e. 100 variants per project, and 600 variants in total. Each variant is checked against all the filters described above, and the pass/fail occurrences are reported.
IV-D Methodology for RQ2: Assessing Variants Uniqueness
To answer RQ2, we assess the diversity of the variants both statically and dynamically.
Diversity at source code level can be removed through the transformations performed at different stages of compilation [35]. We measure how often this occurs, collecting the proportion of variants that preserve uniqueness after compilation with different optimization flags. We compare the unique SHA-256 hashes of the produced machine code. This experiment is performed on the same dataset of C functions as in RQ1 (IV-B). We perform this comparison only for variants that have been proven functionally equivalent previously by the filters applied in IV-C. We use each of four optimization flags -O0, -O1, -O2, -O3 to rebuild the function’s machine code, and compare the SHA256 hash of the produced binaries, to see whether some compilation flags undo specific variants (i.e. make the binary statically identical to the original).
Then, we execute all statically diverse variants, to measure the diversity of observable behaviors among them. We define diverse observable behaviors per the CPU instruction trace. At runtime, we record all CPU instructions executed by the variant. To achieve this, we rely on Intel’s Pin instrumentation tool [36]. We then aggregate and compare the instructions executed by each variant in two ways. First, we compare the CPU instruction set used by the variants and compare it against the set of the original CPU instruction set. This is measured using the Jaccard similarity coefficient [37]. Second, we compare the amount of CPU instructions executed by the variants in terms of absolute numbers.
IV-E Methodology for RQ3: Validating the Ability to Mitigate Miscompilations
As described in subsection III-A, our work specifically focuses on hardening against miscompilation errors produced by bugs in compilers, which is a very dangerous and tricky class of bugs. In RQ3, we consider miscompilation bugs from the Clang compiler. We assess the effectiveness of Galápagos in automatically generating N-Version programs which can mitigate these bugs. We look at all bugs reported in Table III, coming from different versions of Clang and optimization stages. All are unresolved in the repository’s issue tracker at the time of experimentation. For each bug, we create two test programs: one with a function containing a minimal reproducible example for said bug, and one with an N-Version implementation of the same function. We always run Galápagos with the minimal function as the reference broken run. We consider Galápagos to be successful if the N-Version implementation of the reference forces a crash instead of producing the incorrect output due to the miscompilation bug.
IV-F Methodology for RQ4: Measuring Performance Overhead
To measure the performance overhead of N-Version functions generated by Galápagos, we use the clock_gettime(3) API. To get an accurate comparison, we time both the original single-version and the N-Version functions over 2 000 000 executions. Execution and measurements are done in an Intel i7-11800H system, with 16 GB of RAM.
V Experimental Results
In this section, we present and discuss the results of our experiments. In subsection V-A we look at the ability of Galápagos to generate variants that are equivalent to the reference (RQ1). In subsection V-B and subsection V-C, we assess the static and dynamic diversity of the variants, to answer RQ2. We conclude with subsection V-D, where we answer RQ3 with empirical evidence about the N-variants design’s ability to mitigate miscompilation errors.
V-A Validating Variants
Project Function Name Same-language diversification (C) Cross-language diversification (C to Go)
Comp. Isol. Comp. Proj. Test Verif. Comp. Isol. Comp. Proj. Test. Verif.
alaw_to_s16 10 10 10 1 10 10 10 6
iec958_parity 10 8 8 8 10 10 10 10
add 10 9 9 9 10 1 1 1
ulaw_to_s16 10 10 10 9 10 10 10 10
alsa-lib val_seg 10 10 10 2 9 1 1 0
flac_get_max_frame_size 10 10 9 4 10 10 10 0
mix 10 10 10 6 10 10 10 0
vp5_adjust 10 10 10 0 6 0 0 0
weight 10 10 10 10 10 10 10 0
ffmpeg int_sin 10 10 10 0 10 10 10 0
barrett_reduce 10 10 0 0 10 10 10 10
ctz 10 10 10 9 10 10 10 6
int16_t_negative_mask 9 9 7 7 10 8 6 5
int16_t_nonzero_mask 10 10 7 7 9 9 5 5
libgcrypt montgomery_reduce 10 10 6 4 10 10 8 8
fpr_half 10 10 10 2 10 10 10 1
fpr_lt 9 9 9 0 9 9 9 1
modp_montymul 10 10 10 2 10 10 9 0
modp_norm 10 10 7 0 8 8 3 2
liboqs int16_nonzero_mask 10 10 10 3 10 10 10 8
b64_byte_to_char 10 8 8 0 10 6 6 0
b64_byte_to_urlsafe_char 10 8 8 0 10 9 9 0
b64_char_to_byte 10 7 0 0 9 8 0 0
b64_urlsafe_char_to_byte 10 10 0 0 8 0 0 0
libsodium fBlaMka 10 10 10 9 10 10 10 0
icbrt64 10 10 10 7 9 9 9 0
BitDeinterleave 10 10 10 0 10 10 10 3
BitInterleave 10 10 10 0 1 1 1 1
_booth_recode_w5 10 10 10 2 10 10 10 2
openssl _booth_recode_w7 10 10 10 0 10 0 0 0
avg. 99.3% 96.0% 82.6% 32.3% 92.6% 76.3% 69.0% 24.3%
TABLE IV: Count of variant compliance to each validation filter: Compiled in isolation (Comp. Isol.), compiled within project (Comp. Proj.), test suite success (Test), and verified equal (Verif.).
To answer RQ1, we analyze the extent to which the generated variants pass the different validation filters. The numbers are presented in Table IV.
First, we discuss the results for the generation of variants, in the same language. On average, 99,3% (298/300) of the variants are compiled correctly in isolation, and, when inserted in the project, 96,0% (288/300) still successfully compile. Compilation in isolation fails in 2 out of 300 cases: one case because of a syntax error, and the second because the variant calls a function from a library that has not been imported. Compilation within the project fails in 10 out of 298 remaining cases, all of them because the variant’s IR code makes use of a function that is not declared in the project’s IR code. For instance, one of the variants of alsa-lib’s add uses a compiler built-in function, and compiles in isolation, shown in Figure 6. Yet, it fails to compile inside the project, as the built-in function is not declared in the IR scope of the reference function.
The average rate of success decreases for test validation and equivalence verification to 82,6% (248/300) and 32,3% (93/300), respectively. Testing fails in 40 out of 288 remaining cases, either because of at least one failed test case or test suite timeout. An example from ffmpeg’s flac_get_max_frame_size is shown in 7, where constant addition to the count is not performed correctly. Finally, equivalence validation fails in 151 out of 248 remaining cases. Here, we uncover the cases where the LLM is able to synthesize variants that successfully run and behave similarly to the original, yet the complete semantics of the reference function are not correctly captured. Figure 8 shows an equivalence validation failure example for one of the variants of function fBlaMka. The figure shows a snippet of alive-tv’s that indicates that for inputs x and y of 131 074 and 51 111 936 respectively, the reference function outputs 13 398 943 041 538, while the variant outputs 7 235 012 610.
Figure 6: Build failure example: Function ‘__builtin_add_overflow‘ is not declared in the original project’s IR.
Figure 7: A logical error in variant (bottom) which causes test failure. The addition to count is not equivalent.
Figure 8: Formal equivalence validation failure: parameters x and y produce different outputs in the reference function and the variant.
For the cross-language configuration, on average, 92,6% (278/300) of the variants are compiled correctly in isolation, yet, when inserted in the project, only 76,3% (229/300) result in compilation success. Compilation in isolation fails in 22 out of 300 cases, because of syntax errors and failure of compile-time checks, e.g. overflow checks. Figure 9 shows an example where the code fails to compile, as it fails to pass the go compiler’s integer overflow check.
Figure 9: Build in isolation failure example. The code fails with an integer overflow error.
Compilation within the project fails in 49 out of 278 remaining cases, with the reasons behind these failures being (1) the use of IR components in the variant which are not declared in the project’s IR code, and; (2) IR type mismatches. At this stage, an interesting case is openssl’s _booth_recode_w7 function. This function fails to compile within the project because the variants have a modified signature: The LLM added a suffix for each variant, disregarding the prompt’s explicit instructions to keep the original signature.
When executing the variants, the average rate of success drops for test validation and equivalence verification to 69,0% (207/300) and 24,3% (73/300), respectively. Testing fails in 22 out of 229 remaining cases, either because of at least one failed test case or test suite timeout. The acquired errors are logical ones, such as the one displayed in Figure 7, with no major LLM hallucinations detected in the variants generated in our experiments. This is, no nonsensical, or completely off-the-mark solutions. Finally, equivalence validation fails in 110 out of 207 remaining cases, where the LLM misses a part of the semantics of the reference function.
It is worth noting that the acceptability of variants is close for both same-language and cross-language configurations, especially when contrasting the final number of variants that are proved to be equivalent.
Furthermore, for most functions (21/30), both configurations generate a similar amount of equivalent variants. Both same- and cross-language seem to excel or face difficulties for these functions: libsodium’s conversion functions are a clear example of this, consisting of a single line with several chained operations; they appear to confuse the LLM, regardless of the diversification configuration. The remaining functions (9/30) show disparate diversification results across configurations: while for some functions the same-language configuration performs better (5/9), for others the cross-language configuration produces more equivalent variants (4/9).
We now analyze the distribution of equivalent variants per function. In the same-language configuration, for the bottom quartile, 0 equivalent variant were found; for the median, 2 equivalent variants were found; and the top quartile, 7 equivalent variants were found. In the cross-language configuration, for the bottom quartile 0 equivalent variant were found; for the median 1, equivalent variant was found; and for the top quartile, 5 equivalent variants. We can observe that both distributions are skewed towards zero, this means that non-equivalent variants are more likely to be generated.
However, it is notable that for most (23/30) of the functions, at least 1 equivalent variant was generated. We consider this to be a promising result that shows the feasibility of the Galápagos concept. We believe this can be further improved by using more advanced models and sophisticated prompting techniques.
{mdframed}
[style=mpdframe, frametitle=Answer to RQ1] Large language models successfully generate variants given a reference function. Out of 600 function variants, 170 were verifiably equivalent, 97 of which were created from same-language source code and 73 from a source code in a different language. This signals compelling applicability of creating automated cross-language N-Version functions with strong formal guarantees.
V-B Static Uniqueness
Table V shows the number of variants which compile to LLVM-unique IR code, and produce a unique machine code with different optimization flags. The column "Total eq." shows the total number of equivalent variants found, as displayed in Table IV. The column "IR" shows the number of these variants that are unique IR files. We obtain this number by comparing the "Total eq." variants’ hashes at the IR level. Subsequent columns, "-O0" to "-O3", show the number of unique machine code produced by the compiler with each optimization configuration.
For the same-language configuration, 91 of the 93 equivalent variants are different from each other at the IR level. These 91 variants are generated from 18 out of the 30 reference functions. The compilation with "-O0" preserves 80 unique variants. Meanwhile, the number of unique variants is reduced to 72 after "-O3" optimization, representing 74.22% of all equivalent variants. The optimizations observed in the functions are performed at the "-O1" level, and are concerned with inlining, loop invariant code motion, and global value numbering.
For the cross-language configuration, a total of 36 unique variants were found at the IR level, among the 73 equivalent variants.
These 31 variants correspond to 16 out of 30 reference functions. On compilation, the number of unique variants is reduced to 21 after "-O3" optimization, representing 28.76% of all equivalent variants.
A key result for RQ2 is the fact that the introduced diversity is generally well-preserved across optimization stages: from all 127 unique variants at the IR level, a total of 93 (73.22%) are still unique after applying all optimization passes.
In total, at least 1 unique, equivalent variant was found for 24 out of 30 reference functions. Furthermore, we are able to find at least one behaviorally unique variant for each function for which we successfully synthesize at least 1 equivalent variant, . In the case of same-language diversification, the number of unique variants per function is greater than 2 in most cases (15/19). Cross-language diversification, on the other hand, does not perform as well, with only one unique variant for most functions (13/16). We hypothesize that the lower number of unique variants in the cross-language configuration is caused by two reasons: The LLM used in our experiment produces less diverse source code given our experimental parameters, or; the Go-to-IR compiler produces more consistent code across variants.
Project Function Name Same-language diversification (C) Cross-language diversification (C to Go) Totals
Total eq. IR Binary Total eq. IR Binary IR -O3
-O0 -O1 -O2 -O3 -O0 -O1 -O2 -O3
alaw_to_s16 1 1 1 1 1 1 6 4 3 2 2 2 5 3
iec958_parity 4 3 3 3 3 3 4 4 3 2 2 2 7 5
add 9 9 9 9 9 9 1 1 1 1 1 1 10 10
ulaw_to_s16 9 9 8 8 8 8 10 2 1 1 1 1 11 9
alsa-lib val_seg 2 2 2 2 2 2 2 2
flac_get_max_frame_size 4 4 3 3 3 3 4 3
mix 6 6 5 5 5 5 6 5
ffmpeg weight 10 9 8 8 8 8 9 8
barrett_reduce 10 2 1 1 1 1 2 1
ctz 9 9 9 9 9 9 6 5 4 4 4 4 14 13
int16_t_negative_mask 7 5 5 1 1 1 5 3 2 1 1 1 8 2
int16_t_nonzero_mask 7 7 6 4 4 4 5 4 1 1 1 1 11 5
libgcrypt montgomery_reduce 4 4 4 4 4 4 8 1 1 1 1 1 5 5
fpr_half 2 2 2 2 2 2 1 1 1 1 1 1 3 3
fpr_lt 1 1 1 1 1 1 1 1
modp_montymul 2 2 2 2 2 2 2 2
modp_norm 2 1 1 1 1 1 1 1
liboqs int16_nonzero_mask 3 3 2 1 1 1 8 3 1 1 1 1 6 2
libsodium fBlaMka 9 9 6 6 6 6 9 6
icbrt64 7 5 3 2 2 2 5 2
BitDeinterleave 3 1 1 1 1 1 1 1
BitInterleave 1 1 1 1 1 1 1 1
openssl _booth_recode_w5 2 2 2 2 2 2 2 2 1 1 1 1 4 3
Total 97 91 80 72 72 72 73 36 24 21 21 21 127 93
TABLE V: Variant uniqueness. Each column for both configurations shows how many unique variants are obtained at the IR level, and at the machine code level after various optimizations. Functions with no equivalent variants are omitted. The two rightmost columns display the total unique variants for both configurations, both at the IR level and at the machine code level, after building with the -O3 optimization flag.
V-C Dynamic Uniqueness
Now, we execute the diverse, equivalent variants generated by Galápagos. On average, same-language variants’ instruction sets have a Jaccard similarity coefficient of 0.805 with respect to the reference functions’ instruction sets; while the same metric for cross-language variants’ instruction sets is 0.693. In this measurement, the Jaccard similarity coefficient indicates overlap in observable internal behavior. Hence, cross-language variants are strictly better at maximizing diversity at runtime.
Figure 10 shows plots with CPU instruction counts of 6 of the dataset’s functions, one function per project. Each sub-figure shows 3 overlapping areas, aggregating the instruction counts of the original function, an equivalent same-language variant, and an equivalent cross-language variant. In detail, each figure is a radial plot, where the axes represent each of the CPU instructions executed by the function variants. Each colored area aggregates the instruction count of a different function implementation. This means that the less overlap in the figures, the more diverse the variants are in terms of the number and type of executed CPU instructions.
On these plots, we observe that the cross-language variants consistently show more diverse behavior in terms of CPU instructions. For instance, the sub-figure of function _booth_recode_w5 shows an almost complete overlap of the original implementation and the same-language variant’s areas, while the cross-language variant area clearly displays a different set of instructions and count of the overlapping instructions. Similarly, in the case of function ctz, the cross-language variant executes 3 instructions that are executed neither in the reference function nor in the same-language variant, while also avoiding 3 instructions from the original execution set. This difference in instruction subsets follows naturally from using different compiler front-ends for producing the variants’ IR code.
{mdframed}
[style=mpdframe, frametitle=Answer to RQ2] Our experiments show that Galápagos is able to create function variants that are diverse to a large extent. In total, for 166 unique variants across the 30 reference functions’ source code, 126 (75.9%) were still unique after all compilation and optimization passes. This is a very relevant finding, since it means that the approach is valid regardless of optimization requirements.
After executing the variants, we observe diverse internal behavior at runtime, evidenced by different sequences of CPU instructions. This finding is relevant, as it means that the introduced diversity at the machine code level goes beyond minor changes in the binary. It is aligned with the core assumption of N-Version programming, that the fault-tolerance increases with runtime diversity.
Figure 10: Dynamic uniqueness of variants. Each radial plot shows the count of CPU instructions executed by the original function, superimposed with the instruction counts of the same-language (C) and cross-language (Go) variants.
V-D Mitigating Miscompilations
Table VI shows the results of mitigating the Clang miscompilation bugs introduced in Table III with Galápagos. First, we observe that Galápagos successfully generates multiple equivalent variants for each function that triggers each miscompilation bug. Second, Galápagos is able to assemble N-Version binaries for these functions, which caused the execution to crash, instead of producing a wrong output. We consider Galápagos effective in mitigating the three proposed bugs.
Bug Result Configuration #Eq. variants #
M1 Mitigated Same-language 10
M2 Mitigated Same-language 10
M3 Mitigated Same-language 2
TABLE VI: Galápagos’ miscompilation mitigation results.
For bug M1, Galápagos produces 10 equivalent variants in the same-language configuration. It then assembles an 11-Version implementation of the function which triggers the miscompilation fault. The key result here is in the ability of the 11-Version function to let developers know that something wrong has happened. The original function compiles and, when it executes, produces a value. However, the value is wrong because of the miscompilation bug, but there is no warning or check to let the developers know about this wrong value. On the other hand, the 11-Version function compiles, and when it executes, it crashes because the results returned by each version are not consistent. In this case, the program does not silently introduce a wrong value and the developer can act upon this. Consider the control flow graphs in Figure 11. The graph on the left represents the execution of a single-version implementation of the M1 function. The miscompiled code could result in the program silently choosing the wrong execution path, with potentially dangerous results. The graph on the right represents the execution of an N-Version implementation of the M1 function. The miscompiled code executes and returns silently; however, if at least any of the other 10 versions returns a distinct value, the program will exit, thus avoiding wrong execution and informing the program operator that something wrong has occurred. This behavior is similar for the N-Version implementations that Galápagos produced for M2 and M3, with N values of 11 and 3, respectively.
Upon manual inspection of the execution of the resulting binaries, we find that none of the equivalent variants are affected by the miscompilation bugs, only the original function is. As a result, the program termination is triggered by the inconsistent return values between the original function and the return values of one of the corresponding variants.
Figure 11: Control flow graphs for M1. Left: single-version implementation. Right: N-Version implementation.
Figure 12: Disassembled machine code for the hardened M2 function.
Figure 12 shows an excerpt of the machine code produced by Galápagos, when attempting mitigation of M2. The figure shows the calls to the first two versions of the resulting 11-Version implementation: <version_1> is the original function, which is miscompiled; and <version_2> is one of the variants proven to be equivalent at the IR level. In this example, <version_1> returns -1, while <version_2> returns 0. Later in the execution, the return values are compared, and since these are different, the execution is redirected to the ud2 instruction, triggering a crash.
{mdframed}
[style=mpdframe,frametitle=Answer to RQ3] Galápagos is able to generate N-Version programs that mitigate the considered real-world miscompilation bugs. Our results show that equivalent function variants produced by LLMs can be employed to harden critical sections of a given program by assembling N-Version functions.
V-E Performance Overhead
Table VII shows a summary of the overhead introduced by the N-Version functions when compared to the reference functions. The column "Baseline Runtime" shows the measurement taken for each function as described in subsection IV-F. For both same-language and cross-language configurations, "N" represents the number of versions used in the maximally diverse N-version function; "N-Version runtime" shows the execution time for the N-version function; and finally "Normalized Overhead" shows the average overhead for a N-version function with one single-variant (i.e. a 2-version function).
The average normalized overhead is expected to be 2x for a 2-version function, which is what we observe on the last average row. The observed maximum and minimum overheads are 3.75x and 1.04x respectively. The minimum value can be considered an outlier, and it is introduced by a loop-unroll compiler optimization in the Go variant of a C function. We note that it is also possible to design parallel execution harnesses in time-critical contexts.
{mdframed}
[style=mpdframe, frametitle=Answer to RQ4] Our experiments show that the N-Version functions generated by Galápagos introduce the expected overhead. After executing the N-Version functions, we observe that the runtime is increased by 2x for a 2-version function, and grows linearly by adding more variants to the assembly. This experiment demonstrates that the harness overhead overhead is negligible.
Function Name Baseline runtime ($\mu$s) Same-language diversification (C) Cross-language diversification (C+Go)
N N-Version runtime ($\mu$s) Normalized Overhead N N-Version runtime ($\mu$s) Normalized Overhead
alaw_to_s16 4 606 2 17 256 3.75x 7 38 363 2.22x
iec958_parity 83 623 5 269 676 1.56x 5 95 393 1.04x
add 2 486 10 36 362 2.51x 2 5 879 2.36x
ulaw_to_s16 4 369 10 54 707 2.28x 11 52 961 2.11x
val_seg 4 163 3 12 544 2.01x
flac_get_max_frame_size 6 222 5 30 622 1.98x
mix 9 067 7 60 226 1.94x
weight 6 256 11 59 428 1.85x
icbrt64 64 412 8 700 098 2.41x
BitDeinterleave 1 745 4 7 309 2.06x
BitInterleave 1 848 2 4 293 2.32x
_booth_recode_w5 4 599 3 13 757 2.0x 3 11 070 1.7x
fpr_half 2 895 3 12 139 2.6x 2 5 833 2.01x
fpr_lt 5 615 2 13 796 2.46x
modp_montymul 5 543 3 20 754 2.37x
modp_norm 2 793 3 8 210 1.97x
int16_nonzero_mask 3 872 4 15 601 2.01x 9 25 391 1.69x
barrett_reduce 3 095 11 67 948 3.1x
ctz 212 299 10 866 981 1.34x 7 655 523 1.35x
int16_t_negative_mask 2 754 8 25 538 2.18x 6 16 344 1.99x
int16_t_nonzero_mask 4 060 8 27 467 1.82x 6 18 922 1.73x
montgomery_reduce 2 472 5 18 241 2.59x 9 17 304 1.75x
fBlaMka 3 356 10 38 936 2.18x
Avg. Overhead 2.18x 1.99x
TABLE VII: Runtime measurement of N-Version functions.
VI Discussion
In this section, we discuss Galápagos in the larger context of software development, as well as the threats to the validity of our experiments.
VI-A Integration into Existing Software Development Practices
The integration of Galápagos into an existing software development project requires: (1) identifying critical sections of code with the highest reliability standards, and; (2) isolating them into pure functions. Developers would need to integrate Galápagos in the build pipeline for those functions. The pipeline would automatically generate variants, validate their correctness, and insert the N-Version functions into the final build. At deployment time, the N-version functions would terminate execution in case of inconsistencies detected by diversification. This is valid in IoT [38]. We are also aware that companies providing security critical software for the military have such build pipelines with other kinds of diversification.
VI-B Threats to Validity
Threats to internal validity: In our experiments, we use LLVM as the intermediate representation, and alive-tv as the equivalence checker. Therefore, the experimental setting inherits any of their limitations. For instance, some of the variants reported as non-equivalent were not proven as such, but rather the tool failed to process them because of unsupported LLVM instructions.
Also, the inherent non-determinism of LLMs can affect our results. This means that the proportion of equivalent variants can be different in an experimental reproduction. However, non-determinism is key in the design of Galápagos, as it is needed in the diversification pass to generate distinct variants.
Threats to external validity: We identify two dimensions where the obtained results can be generalized for Galápagos’ design. First, our experiments focus on two language pairs: C-to-C and C-to-Go. We argue that similar results can be achieved for other language pairs, as long as these are IR-compatible, and a corresponding IR equivalence checking tool exists. Second, our experiments only consider a single LLM as a source of diversity. We argue that similar results can be achieved using different LLMs, as these have been shown to have similar capabilities [18]. We acknowledge that the generalizability of the results, and the claims of suitability of the design can be strengthened by performing experiments with different language pairs, IRs, and generative models.
VII Related Work
VII-A Automated Variant Synthesis
Singh et al. [39] propose the automatic creation of language-diverse, formally verified program variants from a high-level specification. They focus on fault-tolerance of distributed systems by executing each variant following the replicated state machine paradigm. However, no prototype implementation, experimental methodology, or results are provided. Pelofske et al. [18] present a thorough analysis of function variant generation with LLMs. This study is focused on four SHA-1 functions, and provides a deep insight into the diversity and correctness of LLM-generated variant functions. Xu et al. [17] describe a variant generation mechanism based on LLVM-IR obfuscation. Their goal is to create explicitly non-equivalent variants, which are resistant to replication of known tampering mechanisms. While these works aim at automating variant creation, and to different extents, correctness verification, the key novelty of our work is the focus on hardening of binaries via N-Version function execution.
VII-B LLMs code correctness
Pan et al. [40] identify the correctness of LLM-generated code as an important challenge to address. In this vein, Kessel and Atkinson [41] present LASSO, a platform for differential execution of large sets of code versions and tests. Compared to this work Galápagos goes one step further and automates the creation of N-Version programs, in addition to generation and testing of code variants. Du et al. [42] present a benchmark for solutions to programming tasks via LLMs. This benchmark is designed to measure how efficient are the generated solutions in terms of algorithmic complexity. The experiments show that LLMs are able to produce code that solves the tasks while having diverse efficiency scores. Zhang et al. [28] introduce a benchmark to quantify the correctness of code produced by an LLM. The experiments evidence that LLMs have the capability to correctly solve programming tasks to a large extent. However, while the data provided by these works gives a detailed overview of the correctness and efficiency of code produced by LLMs, the evaluations are test-based, and not formally verified against a specification.
VII-C Code Generation and Code Translation
Code generation and translation with LLMs are very active research areas [43, 44]. Yang et al. [22] present a C-to-Rust verified translation tool. It works by using static methods to construct a known equivalent translation, and LLMs to create more idiomatic translations. The latter are then checked for equivalence against the former using an off-the-shelf Rust tool. Bathia et al. [45] describe a tool to produce verified code transpilations. This is achieved by leveraging intermediate representations and formal equivalence checkers. Yang et al. [46] present UniTrans, a framework to generate and test LLM code translation. UniTrans uses the original code to further generate test cases for the resulting translation, and leverage them to iteratively repair the translation in case of failure. While the mentioned works address the issue of validating correctness of LLM-generated code equivalence, their scope is not concerned with software reliability.
VII-D Code Transplantation and Recombination:
As part of Galápagos’ Harnessing pass, IR code from newly generated programs is inserted into the project that is being hardened. This process is similar to the code transplantation mechanisms described in the literature. This process has been studied in the context of automated program repair [47], feature transfer between systems [48, 49, 50, 51], and code clone analysis [52]. The Harnessing pass is also similar to recombination approaches, where different versions of systems are created by manipulating sections of readily available codebases, at lower-than-source levels [53, 54]. While closely related, these works are not concerned with the challenges of the two first passes of Galápagos, namely, automatic diversification and correctness checking.
VII-E Addressing Miscompilations
The works of Eide et al. [26] and Le et al. [55] present approaches to discover miscompilation bugs via stochastic exploration of the program space. With the same goal, Tu et al. [56] describe a mechanism with guided LLM prompt production approach, which allows producing bug-triggering program mutations. While these approaches seek to actively discover miscompilation bugs, Galápagos takes a program hardening, as it does not target compilers directly, but protects critical application code against potential miscompilation.
VIII Conclusion
In this paper, we have introduced the original idea of automated N-Version programming with LLMs. The key idea is to exploit the search power and coding capabilities of LLMs, within the scope of N-Version programming. We presented Galápagos, a system that allows for automatic N-Version programming, with formal guarantees. We evaluated it in four dimensions: variant correctness, code diversity, the ability to counteract miscompilation bugs, and introduced overhead.
Our experiments show that: (1) Galápagos is able to produce code variants with LLMs, that can be proven to be equivalent, even when these variants are written in a different programming language. (2) Galápagos is able to produce code variants which are diverse both on disk and at execution-time; (3) the Galápagos variants can be harnessed to harden critical sections of software against the dangerous class of miscompilation bugs, and; (4) the N-Version functions generated by Galápagos introduce a performance overhead that stays within reasonable and expected limits. Finally, our experiments clearly show that the production of N-Version software can be automated by the combined use of off-the-shelf tools and generative language models.
We identify two future angles for this novel research direction. Galápagos relies on N-of-N voting, which means that all variants must produce the same output for the program to continue to execute. However, previous works in the N-Version programming line have proposed alternative voting mechanisms [57]. Galápagos can be extended to support different voting mechanisms, so that voting can be tailored to the context and specific requirements of the application. For instance, a threshold voting mechanism (k-of-N) would allow Galápagos to provide similar hardening enhancements, while also hiding unwanted behavior, and continue with execution.
Effective prompt engineering is also crucial in guiding LLMs to generate the desired code variants. For instance, future prompts could include the function’s purpose, input and output specifications, constraints, and examples of desired code patterns. A promising direction is to enrich the prompt with measurements of previously generated variants, such as global alignment, Hamming distance, or Levenshtein distance [58, 59, 60].
Finally, we envision LLMs being fine-tuned for diversity. A fine-tuning loop would help produce code with higher degrees of variant diversity and correctness, as it has been for program repair [61]. Fine-tuning could involve adapting pre-trained models to specific diversification tasks or language pairs, in order to enhance their ability to generate diverse and correct code variants.